# references.bib

### v1

@@ Abstract
Many multi-scale systems exhibit stochastic and deterministic dynamics. Capturing these aspects in a compartmental model is challenging. Notably, low occupancy compartments exhibit stochastic dynamics and high occupancy compartments exhibit deterministic dynamics. Failing to account for stochasticity in small populations can produce ‘atto-foxes’, e.g. in the classic Lotka-Volterra ODE system. This can be resolved with discrete stochastic models e.g. continuous time Markov chains (CTMCs). Unfortunately, simulating CTMCs is infeasible for most realistic populations.
We develop a hybrid CTMC-ODE approach, ‘Jump-Switch-Flow’ (JSF), available as a Python package. JSF couples CTMCs and ODEs, and depending on the state of the system, individual compartments ‘switch’ between stochastic ‘jumping’ and deterministic ‘flowing’ representations. Consequently, compartments can reach extinct states, thereby resolving atto-type problems. JSF has the desired behaviours of exact CTMC simulation, but is substantially faster that existing alternatives.
JSF’s value for inference problems, particularly multi-scale problems, is demonstrated with a case-study using a particle filter. With existing data of SARS-CoV-2 infections, we perform a novel analysis to quantify within-host viral clearance. JSF offers a novel approach to compartmental model development and simulation.
@@
Hybrid simulation — mathematical biology — stochastic modelling — viral clearance — extinction
@@
Dynamical systems represent a powerful method for describing the world and have successfully been applied in many fields. However, modelling populations which change in size across multiple scales remains a challenge. Population sizes in biological processes often change over orders of magnitude, For example, the spread of infectious disease (anderson1991infectious), the boom-and-bust of insect populations (ludwig1978qualitative), and the immune response to infections including HIV perelson2002modelling, and influenza virus (baccam2006kinetics).
@@
In ODEs, the state of the system changes continuously. This allows for “atto-fox problems” where populations can shrink to infeasibly small sizes where a real population would have gone extinct (fowler2021atto; mollison1991dependence; lobry2015migrations). In CTMCs, the discrete state changes stochastically. CTMCs offer a more natural description of small populations sizes, however scaling them to large populations often makes them unwieldy. Moreover, a range of powerful techniques can be brought to bear on ODEs, enabling more thorough analysis. This raises a natural question about what to do when compartment occupancy changes across orders of magnitude.
There are exact and approximate algorithms to simulate CTMCs simoni2019stochastic. Exact methods (e.g. Doob-Gillespie, first/next reaction, and rejection based methods) are computationally expensive when state transitions occur at a high rate (sanft2015constant). Approximate methods (e.g. Tau-leaping (gillespie2001approximate) and chemical Langevin equations (rao2003stochastic; cao2006efficient; gillespie2000chemical; gibson2000efficient)) scale to high transition rates but can have unacceptable approximation error. To overcome the limitations of classical approximate methods, over the past two decades, there has been a concerted effort to devise hybrid stochastic-deterministic approaches (simoni2019stochastic).
Partitioning the transitions in CTMC into fast and slow transitions which can be sampled individually and subsequently synchronised is efficient but complex to implement simoni2019stochastic. Jump-diffusion differential equations partition the model compartments into fluid and discrete to construct hybrid switching jump diffusion processes buckwar2011rungekutta; angius2015approximate.
Essential for the real-world utility of any mathematical model of a dynamical system is its ability to be calibrated to data. For deterministic compartment models, standard parameter inference methods such as maximum likelihood and posterior sampling via Markov chain Monte Carlo (MCMC) can readily be applied (ionides2006inference). For stochastic compartment models these inference problems are usually much harder. Simulation based inference has emerged as a way to handle these problems, e.g. particle filter (arulampalam2002tutorial; kitagawa1996monte).

### v2

@@ Abstract
Multi-scale systems often exhibit stochastic and deterministic dynamics. Capturing these aspects in a compartmental model is challenging. Notably, low occupancy compartments exhibit stochastic dynamics and high occupancy compartments exhibit deterministic dynamics. Failing to account for stochasticity in small populations can produce ‘atto-foxes’, e.g. in the Lotka-Volterra ordinary differential equation (ODE) model. This limitation becomes problematic when studying extinction of species or the clearance of infection, but it can be resolved by using discrete stochastic models e.g. continuous time Markov chains (CTMCs). Unfortunately, simulating CTMCs is infeasible for most realistic populations.
We develop a novel mathematical framework, to couple continuous ODEs and discrete CTMCs: ‘Jump-Switch-Flow’ (JSF). In this framework populations can reach extinct states (“absorbing states”), thereby resolving atto-fox-type problems. JSF has the desired behaviours of exact CTMC simulation, but is substantially faster than existing alternatives.
JSF’s utility for simulation-based inference, particularly multi-scale problems, is demonstrated by several case-studies. In a simulation study, we demonstrate how JSF can enable a more nuanced analysis of the efficacy of public health interventions. We also carry out a novel analysis of longitudinal within-host data from SARS-CoV-2 infections to quantify the timing of viral clearance. JSF offers a novel approach to compartmental model development and simulation.
@@
Hybrid simulation — stochastic modelling —viral clearance — extinction — compartmental modelling
@@
Dynamical systems represent a powerful method for describing the world and have successfully been applied in many fields. However, modelling populations which change in size across multiple scales remains a challenge. Population sizes in biological processes often change over orders of magnitude, e.g. the spread of infectious disease (anderson1991infectious), the boom-and-bust of insect populations (ludwig1978qualitative), and the immune response to infections including HIV (perelson2002modelling), and influenza virus (baccam2006kinetics).
@@
In ODEs, the state of the system changes continuously. This can lead to “atto-fox problems” where populations shrink to infeasibly small sizes where a real population would have gone extinct (fowler2021atto; mollison1991dependence; lobry2015migrations). In CTMCs, the discrete state changes stochastically and there may be absorbing states e.g. extinction of a population. CTMCs offer a more natural description of small populations sizes, however applying them to large populations can be challenging. Moreover, a range of powerful techniques can be brought to bear on ODEs, enabling more thorough analysis. This raises a natural question about what to do when compartment occupancy changes across orders of magnitude.
There are exact and approximate algorithms to simulate CTMCs simoni2019stochastic. Exact methods (e.g. Doob-Gillespie, first/next reaction, and rejection based methods) are computationally expensive when state transitions occur at a high rate (sanft2015constant). Approximate methods (e.g. Tau-leaping (gillespie2001approximate) and chemical Langevin equations (rao2003stochastic; cao2006efficient; gillespie2000chemical; gibson2000efficient)) scale to high transition rates but can have unacceptable approximation error. To overcome the limitations of classical approximate methods, over the past two decades, there has been a concerted effort to develop hybrid stochastic-deterministic approaches (simoni2019stochastic).
Some of the approaches to hybrid modelling involve partitioning the transitions in CTMC into fast and slow transitions which can be sampled individually and subsequently synchronised is efficient but complex to implement simoni2019stochastic. Jump-diffusion differential equations partition the model compartments into fluid and discrete to construct hybrid switching jump diffusion processes buckwar2011rungekutta; angius2015approximate. Recently, kynaston2023regime considered extended systems that represent each compartment with both a continuous and a discrete version and allow the for conversion between them.
Essential for the real-world utility of any mathematical model of a dynamical system is its ability to be calibrated to data. For deterministic compartment models, standard parameter inference methods such as maximum likelihood and posterior sampling via Markov chain Monte Carlo (MCMC) can readily be applied (ionides2006inference). For stochastic compartment models these inference problems are usually much harder. Simulation-based inference has emerged as a way to handle these problems, e.g. particle filter (arulampalam2002tutorial; kitagawa1996monte) and approximate Bayesian computation  (Alahmadi2020ABC).